# Principal stratification with $U$-statistics under principal ignorability

### v3

@@ 2.3 Identification assumptions
Even in experimental studies, the PGCE estimands cannot be identified without additional assumptions because the pair of potential intermediate variables and that of potential final outcomes are never observed simultaneously. To identify the PGCE from the observed data, we require the following assumptions.
@@
Assumption 1 rules out confounding between the treatment assignment and intermediate as well as final outcomes conditional on observed pre-treatment covariates. This assumption holds by design in randomized experiments. For observational studies, Assumption 1 can be satisfied if all confounders affecting the treatment assignment, the intermediate variable, and the final outcome are measured and included in the measured set of pre-treatment covariates, $X$.
@@
Assumption 3 (Mean principal ignorability)
@@
for (a) $(s_{1},s_{2})\neq(s_{1}^{\prime},s_{2}^{\prime})\in\mathcal{S}_{10}=\{(10,10),(11,10),(10,00),(11,00)\}$; (b) $(s_{1},s_{2})\neq(s_{1}^{\prime},s_{2}^{\prime})\in\mathcal{S}_{11}=\{(11,11),(10,11)\}$; (c) $(s_{1},s_{2})\neq(s_{1}^{\prime},s_{2}^{\prime})\in\mathcal{S}_{00}=\{(00,00),(00,10)\}$.
@@
Assumption 3 rules out unmeasured confounding between the principal strata membership and the final outcome, ensuring that the PGCE components can be directly estimated from the observed data. For example, under Assumptions 1-3, the conditional expectations in group $\mathcal{S}_{10}$ are all equal to $\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}|Z_{1}=1,D_{1}=1,Z_{2}=0,D_{2}=0,X_{1},X_{2}]$. In other words, within the observed data stratum defined by $(Z_{1}=1,D_{1}=1,Z_{2}=0,D_{2}=0)$, conditional on pre-treatment covariates, the combinations of possible latent principal strata, $(S_{1}=10,S_{2}=10)$, $(S_{1}=11,S_{2}=10)$, $(S_{1}=10,S_{2}=00)$, and $(S_{1}=11,S_{2}=00)$, are exchangeable with respect to the expected contrast. Similarly, the conditional expected contrasts within group $\mathcal{S}_{11}$ and $\mathcal{S}_{00}$ are all equal to estimable conditional expectations, $\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}|Z_{1}=1,D_{1}=1,Z_{2}=0,D_{2}=1,X_{1},X_{2}]$ and $\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}|Z_{1}=1,D_{1}=0,Z_{2}=0,D_{2}=0,X_{1},X_{2}]$, respectively. That is, with Assumption 3, each one of the subpopulations, $(Z_{1}=1,D_{1}=1,Z_{2}=0,D_{2}=0)$, $(Z_{1}=1,D_{1}=1,Z_{2}=0,D_{2}=1)$, and $(Z_{1}=1,D_{1}=0,Z_{2}=0,D_{2}=0)$, can be viewed as mixtures of its corresponding subgroups defined by observed covariates only.
Remark 1
[…]
Assumptions 2 and 3 are instrumental in achieving nonparametric identification of the PGCE; however, their applicability, in general, cannot be directly tested using observed data alone without prior knowledge. Although our primary focus is identifying and efficiently estimating the PGCE under these two structural assumptions, we provide a confounding function approach for sensitivity analyses under assumed violations of these two structural assumptions in Section S7 of the Online Supplement.
@@ 3.1 Identification of the PGCE estimands
To obtain nonparametric identification formulas for the PGCE estimands, we start by defining components of the observed data distribution, i.e., the treatment probability, the principal score, and the mean of the outcome contrast conditional on covariates (referred to as pairwise outcome mean hereafter for brevity). We consider both the observational and experimental studies and define the treatment probability or the propensity score, by $\pi(X)=\mathbb{P}(Z=1|X)$, which is the probability of a participant receiving the treatment conditional on pre-treatment covariates (Rosenbaum and Rubin, 1983). In addition, the principal score, $e_{s}(X)=\mathbb{P}(S=s|X)$, is defined as the conditional probability that a participant belongs to the principal stratum $s$ given the covariates (Ding and Lu, 2016), for $s\in\{10,00,11\}$ under Assumption 2. Furthermore, we use averages over the distribution of covariates, $\pi=\mathbb{E}\{\pi(X)\}=\mathbb{P}(Z=1)$ and $e_{s}=\mathbb{E}\{e_{s}(X)\}=\mathbb{P}(S=s)$, to denote population fractions of receiving treatment $z$ or belonging to stratum $s$, respectively. Lastly, we define the pairwise outcome mean within the observed stratum $(Z_{1}=z_{1},D_{1}=d_{1},Z_{2}=z_{2},D_{2}=d_{2})$, as
$\mu_{z_{1}d_{1}z_{2}d_{2}}(X_{1},X_{2})=\mathbb{E}\{h(Y_{1},Y_{2})|Z_{1}=z_{1},D_{1}=d_{1},Z_{2}=z_{2},D_{2}=d_{2},X_{1},X_{2}\},$
for $z_{1},z_{2},d_{1},d_{2}\in\{0,1\}$. Below, we provide some examples of the pairwise outcome mean models for specific contrast functions.
Example 3 (Pairwise mean function for average treatment effect)
For studying the average treatment effect, the contrast function is specified as $h(u,v)=u-v$. If the potential outcomes are continuous, a Gaussian regression model is often assumed for the potential outcomes, i.e., $Y(z)|D,X\sim\mathcal{N}(f_{z}(D,X),\sigma^{2})$. Here, $f_{z}(D,X)$ is the mean function of $D$ and $X$ given treatment status $Z=z$. We then have the pairwise outcome mean function given by
$\mu_{z_{1}d_{1}z_{2}d_{2}}(X_{1},X_{2})=f_{z_{1}}(d_{1},X_{1})-f_{z_{2}}(d_{2},X_{2}).$ (7)
Example 4 (Pairwise mean function for probability index)
For studying the probability index estimand, the contrast function is specified as $h(u,v)=\mathbbm{1}(u\geq v)$. If the potential outcomes are continuous, a Gaussian regression model can be considered to represent the mean potential outcomes as in Example 3. In this case, following Wolfe and Hogg (1971), one can derive the pairwise outcome mean as
@@
Consider the win estimands developed in Example 2, where the contrast function is given by the vector $h(u,v)=\{\mathbbm{1}(u>v),\mathbbm{1}(u<v)\}$ (the win estimands are then up to transformations of the two-dimensional estimand). For ordinal outcomes, if we specify the proportional odds or ordinal logistic regression model for $Y(z)$ conditional on $D$ and $X$, Mao (2017) has shown that the resulting pairwise outcome mean functions are given by the two-dimensional vector
@@
=$\quad=\left\{\sum_{q=1}^{Q}\sum_{q^{\prime}=1}^{q-1}p_{z_{1}q}(d_{1},X_{1})p_{z_{2}q^{\prime}}(d_{2},X_{2}),\sum_{q=1}^{Q}\sum_{q^{\prime}=q+1}^{Q}p_{z_{1}q}(d_{1},X_{1})p_{z_{2}q^{\prime}}(d_{2},X_{2})\right\},$
@@
where $p_{zq}(D,X)$ is the conditional probability of $Y(z)=q\in\{1,\ldots,Q\}$ given $D$ and $X$, $Q$ is the largest category of the ordinal outcome.
For estimating the principal scores, we first define the conditional probability, $p_{z}(X)=\mathbb{P}(D=1|Z=z,X)$, with population fraction $p_{z}=\mathbb{E}\{p_{z}(X)\}=\mathbb{P}(D=1|Z=z)$. Then under Assumption 2, the principal scores can be expressed as (Jiang et al., 2022)
$e_{10}(X)=p_{1}(X)-p_{0}(X),~{}~{}~{}e_{00}(X)=1-p_{1}(X),~{}~{}~{}e_{11}(X)=p_{0}(X).$
This is because, with monotonicity, the observed stratum $(Z=0,D=1)$ only contains stratum $S=11$, and observed stratum $(Z=1,D=0)$ only contains stratum $S=00$. The remaining principal score for stratum $S=10$ can thus be obtained by subtracting $e_{11}(X)$ and $e_{00}(X)$ from one. The population fractions of the three principal strata can be similarly connected as $e_{10}=p_{1}-p_{0}$, $e_{00}=1-p_{1}$, and $e_{11}=p_{0}$. We will refer to $\{p_{0}(X),p_{1}(X)\}$ as principal scores due to the one-to-one mapping to $\{e_{10}(X),e_{00}(X),e_{11}(X)\}$. These elements then prepare us to state the set of nonparametric identification results in the following Theorem 1.
Theorem 1 (Nonparametric identification)
Suppose Assumptions 1-3 hold, $e_{s}>0$ for $s\in\{10,00,11\}$, and $10<\pi(X)<$ 1 for all $X\in\mathcal{X}$, the support of covariates. Then, based on different components of the observed data distribution, the following identification formulas for each PGCE can be obtained:
  • (a)
    Based on the treatment probability and the principal score,
    $\tau_{h}^{10}$ $=\mathbb{E}\left[\left\{\frac{e_{10}(X_{1})}{p_{1}-p_{0}}\frac{D_{1}}{p_{1}(X_{1})}\frac{Z_{1}}{\pi(X_{1})}\right\}\left\{\frac{e_{10}(X_{2})}{p_{1}-p_{0}}\frac{1-D_{2}}{1-p_{0}(X_{2})}\frac{1-Z_{2}}{1-\pi(X_{2})}\right\}h(Y_{1},Y_{2})\right],$
    $\tau_{h}^{00}$ $=\mathbb{E}\left[\left\{\frac{1-D_{1}}{1-p_{1}}\frac{Z_{1}}{\pi(X_{1})}\right\}\left\{\frac{e_{00}(X_{2})}{1-p_{1}}\frac{1-D_{2}}{1-p_{0}(X_{2})}\frac{1-Z_{2}}{1-\pi(X_{2})}\right\}h(Y_{1},Y_{2})\right],$
    $\tau_{h}^{11}$ $=\mathbb{E}\left[\left\{\frac{e_{11}(X_{1})}{p_{0}}\frac{D_{1}}{p_{1}(X_{1})}\frac{Z_{1}}{\pi(X_{1})}\right\}\left\{\frac{D_{2}}{p_{0}}\frac{1-Z_{2}}{1-\pi(X_{2})}\right\}h(Y_{1},Y_{2})\right];$
  • (b)
    Based on the treatment probability and the pairwise outcome mean,
    $\tau_{h}^{10}$ $=\mathbb{E}\left(\left[\frac{D_{1}Z_{1}/\pi(X_{1})-D_{1}(1-Z_{1})/\{1-\pi(X_{1})\}}{p_{1}-p_{0}}\right]\right.$
    ×$\qquad\qquad\times\left.\left[\frac{D_{2}Z_{2}/\pi(X_{2})-D_{2}(1-Z_{2})/\{1-\pi(X_{2})\}}{p_{1}-p_{0}}\right]\mu_{1100}(X_{1},X_{2})\right),$
    $\tau_{h}^{00}$ $=\mathbb{E}\left[\left\{\frac{1-D_{1}Z_{1}/\pi(X_{1})}{1-p_{1}}\right\}\left\{\frac{1-D_{2}Z_{2}/\pi(X_{2})}{1-p_{1}}\right\}\mu_{1000}(X_{1},X_{2})\right],$
    $\tau_{h}^{11}$ $=\mathbb{E}\left(\left[\frac{D_{1}(1-Z_{1})/\{1-\pi(X_{1})\}}{p_{0}}\right]\left[\frac{D_{2}(1-Z_{2})/\{1-\pi(X_{2})\}}{p_{0}}\right]\mu_{1101}(X_{1},X_{2})\right);$
  • (c)
    Based on the principal score and the pairwise outcome mean,
    $\tau_{h}^{10}$ $=\mathbb{E}\left[\left\{\frac{p_{1}(X_{1})-p_{0}(X_{1})}{p_{1}-p_{0}}\right\}\left\{\frac{p_{1}(X_{2})-p_{0}(X_{2})}{p_{1}-p_{0}}\right\}\mu_{1100}(X_{1},X_{2})\right],$
    $\tau_{h}^{00}$ $=\mathbb{E}\left[\left\{\frac{1-p_{1}(X_{1})}{1-p_{1}}\right\}\left\{\frac{1-p_{1}(X_{2})}{1-p_{1}}\right\}\mu_{1000}(X_{1},X_{2})\right],$
    $\tau_{h}^{11}$ $=\mathbb{E}\left[\left\{\frac{p_{0}(X_{1})}{p_{0}}\right\}\left\{\frac{p_{0}(X_{2})}{p_{0}}\right\}\mu_{1101}(X_{1},X_{2})\right].$
We use $\tau_{h}^{10}$ as an example to elaborate on the intuitions behind Theorem 1, and the identification of the other two PGCEs follows the same strategy. In Theorem 1(a), $\tau_{h}^{10}$ is expressed as a doubly inverse probability-weighted average of the observed contrast function; the essential part of the weighting scheme leverages the principal scores to tilt all pairs of $(\mathcal{O}_{i},\mathcal{O}_{j})$ ($i\neq j$) in the observed stratum $(Z_{1}=1,D_{1}=1,Z_{2}=0,D_{2}=0)$ to represent the covariate distribution in the principal stratum $S=10$. Specifically, under Assumptions 1 and 2, we have
$\mathbb{E}\left[\left\{\frac{e_{10}(X_{1})}{p_{1}-p_{0}}\frac{D_{1}}{p_{1}(X_{1})}\frac{Z_{1}}{\pi(X_{1})}\right\}\left\{\frac{e_{10}(X_{2})}{p_{1}-p_{0}}\frac{1-D_{2}}{1-p_{0}(X_{2})}\frac{1-Z_{2}}{1-\pi(X_{2})}\right\}\bigg{\rvert}X_{1},X_{2}\right]=\frac{e_{10}(X_{1})e_{10}(X_{2})}{e_{10}^{2}},$
which is proportional to the conditional probability of each pair belonging to the principal stratum $S=10$, and by Bayes’ Theorem, is equivalent to the density ratio of $(X_{1},X_{2})$ conditional and unconditional on $S_{1}=S_{2}=10$. This ensures the identification of $\tau_{h}^{10}$. In Theorem 1(b), while $\tau_{h}^{10}$ is expressed using the treatment probability and the pairwise outcome mean, it can be shown that the conditional expectation of the weight attached to the pairwise outcome mean, under Assumptions 1 and 2, is,
$\mathbb{E}\left(\left[\frac{D_{1}Z_{1}/\pi(X_{1})-D_{1}(1-Z_{1})/\{1-\pi(X_{1})\}}{p_{1}-p_{0}}\right]\right.$
×$\qquad\times\left.\left[\frac{D_{2}Z_{2}/\pi(X_{2})-D_{2}(1-Z_{2})/\{1-\pi(X_{2})\}}{p_{1}-p_{0}}\right]\bigg{\rvert}X_{1},X_{2}\right)=\frac{e_{10}(X_{1})e_{10}(X_{2})}{e_{10}^{2}}.$
Therefore, $\tau_{h}^{10}$ is identified as the pairwise outcome mean function averaged over the density $f(X_{1},X_{2}|S_{1}=S_{2}=10)/f(X_{1},X_{2})$. Finally, in Theorem 1(c), $\tau_{h}^{10}$ is expressed using the principal score and pairwise outcome mean, where, in comparison with Theorem 1(b), the treatment probability and intermediate outcome, $DZ/\pi(X)-D(1-Z)/\{1-\pi(X)\}$, is replaced by the principal score, $p_{1}(X)-p_{0}(X)$. The conditional expectation of the weight on covariates attached to the pairwise outcome mean, under Assumptions 1 and 2, remains $e_{10}(X_{1})e_{10}(X_{2})/e_{10}^{2}$, and hence $\tau_{h}^{10}$ is identified as an average of the pairwise outcome mean function over the target covariate density.

### v4

@@ 2.3 Identification assumptions
To identify the PGCE from the observed data, we require the following assumptions.
@@
Assumption 1 rules out confounding between the treatment assignment and intermediate variable as well as final outcomes conditional on observed pre-treatment covariates. This assumption holds by design in randomized experiments. For observational studies, Assumption 1 can be satisfied if all confounders (common causes) affecting the treatment assignment and the intermediate variable, and those affecting the treatment assignment and the final outcome are measured and included in $X$.
@@
Assumption 3 (Pairwise mean principal ignorability)
@@
if $(s_{1},s_{2})$ and $(s_{1}^{\prime},s_{2}^{\prime})$ belong to the same set defined by one of the following: (a) $\mathcal{S}_{10}=\{(10,10),(10,00),(11,10),(11,00)\}$; (b) $\mathcal{S}_{11}=\{(10,11),(11,11)\}$; (c) $\mathcal{S}_{00}=\{(00,10),(00,\allowbreak 00)\}$; and (d) $\mathcal{S}_{01}=\{(00,11)\}$.
@@
Intuitively, Assumption 3 rules out unmeasured confounding between principal stratum membership and the final outcome, thereby ensuring that the PGCE estimands can be identified from the observed data. Under Assumptions 1 - 3, for instance, the conditional expectations within group $\mathcal{S}_{10}$ all reduce to $\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}|Z_{1}=1,D_{1}=1,Z_{2}=0,D_{2}=0,X_{1},X_{2}]$. In other words, within the observed stratum $(Z_{1}=1,D_{1}=1,Z_{2}=0,D_{2}=0)$, conditional on pre-treatment covariates, the various combinations of possible latent strata including $(S_{1}=10,S_{2}=10)$, $(S_{1}=11,S_{2}=10)$, $(S_{1}=10,S_{2}=00)$, and $(S_{1}=11,S_{2}=00)$ are exchangeable with respect to the expected contrast. Similarly, the conditional expected contrasts within $\mathcal{S}_{11}$, $\mathcal{S}_{00}$, and $\mathcal{S}_{01}$ are all equal to the estimable conditional expectations, $\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}|Z_{1}=1,D_{1}=1,Z_{2}=0,D_{2}=1,X_{1},X_{2}]$, $\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}|Z_{1}=1,D_{1}=0,Z_{2}=0,D_{2}=0,X_{1},X_{2}]$, and $\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}|Z_{1}=1,D_{1}=0,Z_{2}=0,D_{2}=1,X_{1},X_{2}]$, respectively (see Table 1(a) for a concrete mapping between the observed data strata and the latent principal strata). That is, with Assumption 3, each one of the observed subpopulations, $(Z_{1}=1,D_{1}=1,Z_{2}=0,D_{2}=0)$, $(Z_{1}=1,D_{1}=1,Z_{2}=0,D_{2}=1)$, $(Z_{1}=1,D_{1}=0,Z_{2}=0,D_{2}=0)$, and $(Z_{1}=1,D_{1}=0,Z_{2}=0,D_{2}=1)$, can be viewed as mixtures of its corresponding subgroups defined by observed covariates only (Table 1(b)).
Assumption 3 can be viewed as a generalization of the unitwise mean principal ignorability assumption in Jiang et al. (2022), since it reduces to that form in the linear contrast case h(u,v)=uvh(u,v)=u-v. For nonlinear contrasts, however, expectations cannot be interchanged within the contrast function, and thus there is no clear nesting relationship between the pairwise and unitwise versions of mean principal ignorability. Consider the win estimands under noncompliance (Example 2) with the contrast function $h(u,v)=\mathbbm{1}(u>v)$. Applying (LABEL:eq:pi-equation) to the first two pairs in $\mathcal{S}_{10}$ yields $P\{Y_{1}(1)>Y_{2}(0)|S_{1}=10,S_{2}=10,X_{1}=x_{1},X_{2}=x_{2}\}=P\{Y_{1}(1)>Y_{2}(0)|S_{1}=11,S_{2}=10,X_{1}=x_{1},X_{2}=x_{2}\}$. This assumption states that, conditional on the same pairwise covariate profiles, the probability that a treated individual has a more favorable outcome than a control individual is the same whether the pair consists of two compliers or of one complier and one always-taker. In other words, there do not exist additional unmeasured factors influencing the conditional probability of winning among pairs in $\mathcal{S}_{10}$. This is neither stronger nor weaker than the unitwise mean principal ignorability assumption, which requires, for example, $\mathbb{E}\{Y_{1}(1)|S_{1}=10,X_{1}=x_{1}\}=\mathbb{E}\{Y_{1}(1)|S_{1}=11,X_{1}=x_{1}\}$.
Table 1: Correspondence between latent principal strata pairs, Assumption 3, and observed strata pairs defined by assignment and observed intermediate variable.
(a) The mapping from latent principal strata pairs to the conditional pairwise mean defined in observed strata, through Assumption 3. There are nine latent strata pairs under monotonicity.
Latent strata pairs Conditions in Pairwise mean defined in observed strata
$S_{1}$ $S_{2}$ Assumption 3 $\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}|Z_{1},D_{1},Z_{2},D_{2},X_{1},X_{2}]$
10 10 (a) $\{Z_{1}=1,D_{1}=1\}$ & $\{Z_{2}=0,D_{2}=0\}$
10 $0$ (a) $\{Z_{1}=1,D_{1}=1\}$ & $\{Z_{2}=0,D_{2}=0\}$
11 10 (a) $\{Z_{1}=1,D_{1}=1\}$ & $\{Z_{2}=0,D_{2}=0\}$
11 $0$ (a) $\{Z_{1}=1,D_{1}=1\}$ & $\{Z_{2}=0,D_{2}=0\}$
10 11 (b) $\{Z_{1}=1,D_{1}=1\}$ & $\{Z_{2}=0,D_{2}=1\}$
11 11 (b) $\{Z_{1}=1,D_{1}=1\}$ & $\{Z_{2}=0,D_{2}=1\}$
$0$ 10 (c) $\{Z_{1}=1,D_{1}=0\}$ & $\{Z_{2}=0,D_{2}=0\}$
$0$ $0$ (c) $\{Z_{1}=1,D_{1}=0\}$ & $\{Z_{2}=0,D_{2}=0\}$
$0$ 11 (d) $\{Z_{1}=1,D_{1}=0\}$ & $\{Z_{2}=0,D_{2}=1\}$
(b) The mapping from observed strata pairs (defined by assignment and observed intermediate variable) as a mixture of latent principal strata pairs, through Assumption 3. As the estimand is constructed by contrasting $Y_{1}(1)$ with $Y_{2}(0)$, only four observed strata pairs that satisfy $Z_{1}=1$ and $Z_{2}=0$ are informative for identification.
Observed strata pairs Conditions in Mixture components of
Unit 1 Unit 2 Assumption 3 latent strata pairs
$\{Z_{1}=1,D_{1}=1\}$ $\{Z_{2}=0,D_{2}=0\}$ (a) $(S_{1},S_{2})\in\mathcal{S}_{10}$
$\{Z_{1}=1,D_{1}=1\}$ $\{Z_{2}=0,D_{2}=1\}$ (b) $(S_{1},S_{2})\in\mathcal{S}_{11}$
$\{Z_{1}=1,D_{1}=0\}$ $\{Z_{2}=0,D_{2}=0\}$ (c) $(S_{1},S_{2})\in\mathcal{S}_{00}$
$\{Z_{1}=1,D_{1}=0\}$ $\{Z_{2}=0,D_{2}=1\}$ (d) $(S_{1},S_{2})\in\mathcal{S}_{01}$
Finally, we distinguish Assumptions 1 and 2 from Assumption 3. The first two are stated at the individual level, but since the observed data are assumed to be i.i.d., they extend naturally to pairs. For example, individual-level treatment ignorability implies that $(Z_{1},Z_{2})$ is jointly independent of the potential outcomes and intermediate variables given $(X_{1},X_{2})$. By contrast, Assumption 3 must be formulated at the pairwise level in its minimally sufficient form, because identification with nonlinear contrasts requires linking expectations across observed and latent strata pairs as in Table 1. Assumptions 2 and 3 are not testable from the observed data alone. We therefore complement our main results with sensitivity analyses under violations of these assumptions, provided in Section 4.4 and Section S7 of the Online Supplement.
@@ 3.1 Identification of the PGCE estimands
To obtain nonparametric identification formulas for the PGCE estimands, we start by defining components of the observed data distribution. We consider both the observational and experimental studies and define the treatment propensity score by $\pi(X)=\mathbb{P}(Z=1|X)$, which is the probability of a participant receiving the treatment conditional on pre-treatment covariates (Rosenbaum and Rubin, 1983). In addition, the principal score, $e_{s}(X)=\mathbb{P}(S=s|X)$, is defined as the conditional probability that a participant belongs to the principal stratum $s$ given the covariates (Ding and Lu, 2016), for $s\in\{10,00,11\}$ under Assumption 2. Furthermore, we use $\pi=\mathbb{E}\{\pi(X)\}=\mathbb{P}(Z=1)$ and $e_{s}=\mathbb{E}\{e_{s}(X)\}=\mathbb{P}(S=s)$ to denote population fractions of receiving treatment $z$ or belonging to stratum $s$, respectively. Lastly, we define the pairwise outcome mean within the observed stratum $(Z_{1}=z_{1},D_{1}=d_{1},Z_{2}=z_{2},D_{2}=d_{2})$, as $\mu_{z_{1}d_{1}z_{2}d_{2}}(X_{1},X_{2})=\mathbb{E}\{h(Y_{1},Y_{2})|Z_{1}=z_{1},D_{1}=d_{1},Z_{2}=z_{2},D_{2}=d_{2},X_{1},X_{2}\}$, for $z_{1},z_{2},d_{1},d_{2}\in\{0,1\}$. Below, we provide two examples of the pairwise outcome mean models for specific contrast functions.
Example 3 (Pairwise mean function for probabilistic index)
For studying the probabilistic index estimand, the contrast function is specified as $h(u,v)=\mathbbm{1}(u\geq v)$. If the potential outcomes are continuous, a Gaussian regression model can be considered such that $Y(z)|D,X\sim\mathcal{N}(f_{z}(D,X),\sigma^{2})$. In this case, one can derive the pairwise outcome mean as
@@
Consider the win estimands developed in Example 2, where the contrast function is given by the vector $h(u,v)=\{\mathbbm{1}(u>v),\mathbbm{1}(u<v)\}$. For ordinal outcomes, if we specify the proportional odds or ordinal logistic regression model for $Y(z)$ conditional on $D$ and $X$, Mao (2017) has shown that the resulting pairwise outcome mean functions are given by the two-dimensional vector
@@
$\quad=\left\{\sum_{q=2}^{Q}\sum_{q^{\prime}=1}^{q-1}p_{z_{1}q}(d_{1},X_{1})p_{z_{2}q^{\prime}}(d_{2},X_{2}),\sum_{q=1}^{Q-1}\sum_{q^{\prime}=q+1}^{Q}p_{z_{1}q}(d_{1},X_{1})p_{z_{2}q^{\prime}}(d_{2},X_{2})\right\},$
@@
where $p_{zq}(D,X)$ is the conditional probability of $Y(z)=q\in\{1,\ldots,Q\}$ given $D$ and $X$, and $Q(\geq 2)$ is the largest category of the ordinal outcome.
For estimating the principal scores, we first define the conditional probability, $p_{z}(X)=\mathbb{P}(D=1|Z=z,X)$, with population fraction $p_{z}=\mathbb{E}\{p_{z}(X)\}=\mathbb{P}(D=1|Z=z)$. Then, under Assumption 2, the principal scores can be expressed as $e_{10}(X)=p_{1}(X)-p_{0}(X)$, $e_{00}(X)=1-p_{1}(X)$, and $e_{11}(X)=p_{0}(X)$. The population fractions of the three principal strata can be similarly connected as $e_{10}=p_{1}-p_{0}$, $e_{00}=1-p_{1}$, and $e_{11}=p_{0}$. We will refer to $\{p_{0}(X),p_{1}(X)\}$ as principal scores due to the one-to-one mapping to $\{e_{10}(X),e_{00}(X),e_{11}(X)\}$. Based on these elements, we establish three alternative nonparametric identification formulas for each PGCE estimand, as shown in Theorem S2 of the Online Supplement. Each formula requires specifying only two of the three models—the treatment propensity score, the principal score, and the pairwise outcome mean—rather than all three simultaneously. Furthermore, Theorem S1 of the Online Supplement presents the associated balancing conditions for any pairwise covariate functions. These balancing conditions provide the theoretical foundation for model checking of the treatment propensity and principal score models without peeking at the outcome data.
3.2 Estimators based on nonparametric identification formulas
For point estimation, a common practice for implementing the nonparametric identification formulas is through specifying parametric regression models for nuisance functions, i.e., $\pi(X;\alpha)$, $p_{z}(X;\beta)$ or $e_{s}(X;\beta)$, and $\mu_{z_{1}d_{1}z_{2}d_{2}}(X_{1},X_{2};\gamma)$, where $\alpha$, $\beta$, and $\gamma$ are finite-dimensional nuisance parameters, with their corresponding true values denoted by $\alpha^{*}$, $\beta^{*}$, and $\gamma^{*}$. We write the maximum likelihood estimators for these finite-dimensional parameters as $\widehat{\alpha}$, $\widehat{\beta}$, and $\widehat{\gamma}$. To aid further discussions of statistical properties of estimators for the PGCEs, we introduce notations $\mathcal{M}_{\pi}$, $\mathcal{M}_{p}$, and $\mathcal{M}_{\mu}$ to denote conditions where the treatment propensity, the principal score, and the outcome mean models are correctly specified, respectively. That is, under $\mathcal{M}_{\pi}$, we have $\pi(X;\alpha^{*})=\pi(X)$; under $\mathcal{M}_{p}$, we have $p_{z}(X;\beta^{*})=p_{z}(X)$ and $e_{s}(X;\beta^{*})=e_{s}(X)$ for $z\in\{0,1\}$ and $s\in\{10,00,11\}$; and under $\mathcal{M}_{\mu}$, we have $\mu_{z_{1}d_{1}z_{2}d_{2}}(X_{1},X_{2};\gamma^{*})=\mu_{z_{1}d_{1}z_{2}d_{2}}(X_{1},X_{2})$ for $z_{1},z_{2},d_{1},d_{2}\in\{0,1\}$. Additionally, we use the logic notation ‘+’ to denote conditions where more than one nuisance function is correctly specified, and $\cup$’ to denote the condition where at least one nuisance function is correctly specified.
We can obtain plug-in estimators by replacing components in the identification formulas with their empirical counterparts and expectations with empirical averages. For convenience, we first introduce the following notations: $\mathbb{E}_{n}f(\mathcal{O})=n^{-1}\sum_{i=1}^{n}f(\mathcal{O}_{i})$ for any $f(\mathcal{O})$, and $\mathbb{U}_{n}f(\mathcal{O}_{1},\mathcal{O}_{2})=\binom{n}{2}^{-1}\sum_{i<j}f(\mathcal{O}_{i},\mathcal{O}_{j})$ for any $f(\mathcal{O}_{1},\mathcal{O}_{2})$. Define the following quantity for any function $f(Y,D,X)$:
$\begin{aligned} \psi_{f(Y_{z},D_{z},X)}(\mathcal{O})&=\frac{\mathbbm{1}(Z=z)[f(Y,D,X)-\mathbb{E}\{f(Y,D,X)|X,Z=z\}]}{\mathbb{P}(Z=z|X)}\\ &\qquad+\mathbb{E}\{f(Y,D,X)|X,Z=z\}\end{aligned}.$ (9)
One example of (9) is $f(Y_{z},D_{z},X)=D_{z}$, in which case we obtain the following function:
(10) is useful for the estimation of $p_{z}$, as it leads to a doubly robust estimator once we treat the intermediate variable as the outcome:
[…]
For any general constrast function $h$, the plug-in estimators for the PGCEs take the form of $U$-statistics (van der Vaart, 1998, §12). For example, the identification results in Theorem S2 of the Online Supplement motivate the following estimators for $\tau_{h}^{10}$:
$\widehat{\tau}_{h,\pi+p}^{10}$ $=\mathbb{U}_{n}\left[\left\{\frac{e_{10}(X_{1};\widehat{\beta})}{\widehat{p}_{1}-\widehat{p}_{0}}\frac{D_{1}}{p_{1}(X_{1};\widehat{\beta})}\frac{Z_{1}}{\pi(X_{1};\widehat{\alpha})}\right\}\left\{\frac{e_{10}(X_{2};\widehat{\beta})}{\widehat{p}_{1}-\widehat{p}_{0}}\frac{1-D_{2}}{1-p_{0}(X_{2};\widehat{\beta})}\frac{1-Z_{2}}{1-\pi(X_{2};\widehat{\alpha})}\right\}h(Y_{1},Y_{2})\right],$
[…]
Upon closer investigation, $\widehat{\tau}_{h,\pi+p}^{10}$ is based on the estimated treatment propensity score and the principal score, $\widehat{\tau}_{h,\pi+\mu}^{10}$ is based on the estimated treatment propensity score and the pairwise outcome mean, and $\widehat{\tau}_{h,p+\mu}^{10}$ is based on the estimated principal score and the pairwise outcome mean. For brevity, we will use $\tau_{h}^{10}$ as the running example for exposition in the remainder of the article. Results for $\tau_{h}^{00}$ and $\tau_{h}^{11}$ are analogous and are given in the middle of Section S2 of the Online Supplement. These estimators have strong requirements on model specification to consistently estimate the PGCEs. To reiterate, $\widehat{\tau}_{h,\pi+p}^{s}$ for $s\in\{10,00,11\}$ are consistent under $\mathcal{M}_{\pi+p}$, that is, when the treatment propensity and the principal score are both correctly specified; $\widehat{\tau}_{h,\pi+\mu}^{s}$ for $s\in\{10,00,11\}$ are consistent under $\mathcal{M}_{\pi+\mu}$, since the doubly robust estimator for $p_{z}$ is consistent under $\mathcal{M}_{\pi}\cup\mathcal{M}_{p}$; and $\widehat{\tau}_{h,p+\mu}^{s}$ for $s\in\{10,00,11\}$ are consistent under $\mathcal{M}_{p+\mu}$.